OpenStudy (anonymous):

will fan and medal. Eric is 6 1/6 feet tall and his brother is 5 3/4 feet tall. Eric is how many feet taller than his brother?

OpenStudy (anonymous):

answer choices 5/12 feet 7/12 feet 3/4 feet 1/6 feet

OpenStudy (anonymous):

@iGreen if u wouldn't mind can u help me again

OpenStudy (igreen):

Subtract..first we can convert into improper fractions.. \(6\dfrac{1}{6} \rightarrow \dfrac{37}{6}\) \(5 \dfrac{3}{4} \rightarrow \dfrac{23}{4}\) Now we have to change them to like denominators.. \(\dfrac{37}{6} \rightarrow \dfrac{37 \times 2}{6 \times 2} \rightarrow \dfrac{74}{12}\) \(\dfrac{23}{4} \rightarrow \dfrac{23 \times 3}{4 \times 3} \rightarrow \dfrac{69}{12}\) So we have: \(\dfrac{74}{12} - \dfrac{69}{12} \rightarrow \dfrac{74-69}{12}\)

OpenStudy (anonymous):

so 5/12

OpenStudy (anonymous):

is that right

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thx again

OpenStudy (igreen):

No problem.